Find a polynomial for the perimeter and for the area.
The rectangle shows that the length is n+2 and that the width is n.
The perimeter is ____.
The area is ____.
Simplify your answers but do not factor.

Answer by stanbon(75887)

Find a polynomial for the perimeter and for the area.
The rectangle shows that the length is n+2 and that the width is n.
The perimeter is 2(length+width) = 2(n+2+n) = 2(2n+2) = 4n+4
----
The area is n(n+2) = n^2+2n
============
Cheers,
Stan H.